If the age of father and son is 84.The age of father is three times of son then find the age of father and son.also show in graph.
Answer by stanbon(75887)

If the age of father and son is 84.The age of father is three times of son then find the age of father and son.also show in graph.
-------
Equations:
f + s = 84
f = 3s
----
Substitute for "f" and solve for "s":
3s+s = 84
4s = 84
s = 21 (son's age now)
-----
Solve for "f":
f = 3s
f = 3*21
f = 63 (father's age now)
=============================
cheers,
Stan H.
==========